a block of mass 10kg on a rough horizontal surface is connected to all of mass M2=5kg by a light weight cord over a light weight friction less pully

A force of 120N at angle 53° of the horizontal applied to the block of mass 10kg the cofficent of kinetic friction b/n the surface and the block 0.3 then what is

The magnitude b/n the 2 blocks

The acceleration of the blocks

The tension b/n the 2 blocks


1
Expert's answer
2022-02-11T15:30:53-0500

"m_1\vec a=\vec F+m_1\vec g+\vec N+\vec F_{fr}+m_2\vec g,"

"m_1a=T\cos\alpha-\mu m_1g-m_2g,"

1)

"N=m_1g=98~N,"

2)

"a=\frac F {m_1}\cos\alpha-\mu g-\frac{m_2}{m_1}g=-0.6~\frac m{s^2},"

3)

"T_1=T_2=m_2g=49~N."